OpenStudy (anonymous):

Divide: (7 + 4i) ÷ (2 + 3i) I've been getting crazy numbers.

OpenStudy (anonymous):

2i

OpenStudy (anonymous):

That's not one of my choices.

OpenStudy (anonymous):

what are your choices

OpenStudy (anonymous):

A.) -2+i B.) 2-i C.) -2-i D.) 2=i

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

D.) 2+i

OpenStudy (anonymous):

Answer is B 2-1. You can use this website called Mathway.com and plug in your equations and it gives you the answers you need.

OpenStudy (anonymous):

2-i

OpenStudy (anonymous):

under the tab pre algrebra

OpenStudy (anonymous):

\[\frac{ 7+4i }{ 2+3i}=\frac{ 7+4i }{ 2+3i} \times\frac{ 2-3i }{ 2-3i}=\frac{ \left( 7+4i \right)\left( 2-3i \right) }{ 2^2+3^2 }=\frac{ 14+12+8i-21i }{ 13 }=2-i\]

OpenStudy (anonymous):

Wow. I didn't even think of writing it out that way. Thank you, now I actually can comprehend it. @CarlosGP

OpenStudy (anonymous):

Oh and thank you @whydoihavetosignup1 for showing me a great site!

OpenStudy (anonymous):

welcome. whenever you have a complex number in the denominator, just multyply numerator and denominator by the denominator´s conjugated. like this you will get a real number in the denominator

OpenStudy (anonymous):

your welcome